# Geolocation with Real Human Gameplay Data: A Large-Scale Dataset and Human-Like Reasoning Framework

### v1

@@ Abstract.
Geolocation, the task of identifying an image’s location, requires complex reasoning and is crucial for navigation, monitoring, and cultural preservation. However, current methods often produce coarse, imprecise, and non-interpretable localization. A major challenge lies in the quality and scale of existing geolocation datasets. These datasets are typically small-scale and automatically constructed, leading to noisy data and inconsistent task difficulty, with images that either reveal answers too easily or lack sufficient clues for reliable inference. To address these challenges, we introduce a comprehensive geolocation framework with three key components: GeoComp, a large-scale dataset; GeoCoT, a novel reasoning method; and GeoEval, an evaluation metric, collectively designed to address critical challenges and drive advancements in geolocation research. At the core of this framework is GeoComp (Geolocation Competition Dataset), a large-scale dataset collected from a geolocation game platform involving 740K users over two years. It comprises 25 million entries of metadata and 3 million geo-tagged locations spanning much of the globe, with each location annotated thousands to tens of thousands of times by human users. The dataset offers diverse difficulty levels for detailed analysis and highlights key gaps in current models. Building on this dataset, we propose Geographical Chain-of-Thought (GeoCoT), a novel multi-step reasoning framework designed to enhance the reasoning capabilities of Large Vision Models (LVMs) in geolocation tasks. GeoCoT improves performance by integrating contextual and spatial cues through a multi-step process that mimics human geolocation reasoning. Finally, using the GeoEval metric, we demonstrate that GeoCoT significantly boosts geolocation accuracy by up to 25% while enhancing interpretability. Access dataset  here.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Google-WS-15k  (Clark et al., 2023a) 15k Global Map Service
GMCP (Zamir and Shah, 2014) 105K Local Map Service
StreetCLIP (Haas et al., 2023) 1M Unknown Map Service
@@
Table 1. Comparison of Existing Geolocation Datasets and GeoComp. GeoComp is the first to include real, rich player performance data. “Local” refers to city- or region-specific data, while “Global” spans multiple continents. Darker green shades indicate broader geographic coverage.
@@
Figure 1. The gaming logic of our platform: Two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground truth location.
To address the above obstacles, in this work, we leverage the contributions of hundreds of thousands of geolocation game enthusiasts who provide valuable annotation labels while playing the game. Specifically, we launched a free, public-benefit-oriented online geoguessing website in June 2022 (tuxun.fun), as shown in Figure 1. Users are required to guess the latitude and longitude based on these images, which are sourced from Google Maps, Baidu Maps, and Gaode Maps. The platform offers various game formats, including team contests and solo matches. As of December 17, 2024, this platform has 740,468 users, 3,954,397 locations as unique geolocation tasks, and 25,355,174 human response records. We name the collected dataset GeoComp. This rich and valuable dataset of real human responses enables us to evaluate task difficulty and filter out unreasonable cases. For instance, some tasks are too easy, such as when the name of a shopping mall in a city is clearly visible in the image, enabling most users to answer correctly. On the other hand, some tasks are highly challenging, where only a few users spend considerable time before providing accurate answers.
@@
Our work makes key contributions to geolocation. First, we present GeoComp, a large-scale, human-annotated geolocation dataset with over 3.9 million tasks and 25 million human player annotations, featuring diverse geographic regions, languages, and environmental contexts. These annotations identify high-difficulty geolocation cases and establish benchmarks to guide future advancements. Second, we introduce the Geographical Chain of Thought (GeoCoT) framework, a multi-step reasoning approach that improves geolocation accuracy by leveraging geographical cues like landmarks, environmental features, and spatial relationships. Finally, through comprehensive evaluations involving human assessments and LLM inferences, we show that GeoCoT improves predictive performance by up to 25% while enhancing interpretability.
Figure 2. Spatial distribution of 3,238,919 geo-tagged locations in GeoComp: (a) The global map shows spatial heterogeneity, with dense clusters in more urbanized regions like Europe and Asia, and sparse coverage in areas like Africa and Siberia. (b) The pie chart highlights the proportional geo-tagged locations distribution, led by North America and Asia. (c) Unlike previous datasets like OSV-5M, where a single country (e.g., America) dominates 25% of the data, our dataset is balanced at country level.
@@ 2.1. Image Geolocation Task
Image geolocation refers to determining the corresponding location of a given image, a crucial task in computer vision (Zhu et al., 2023b, a, 2024a), spatial data mining (Zhao et al., 2017, 2016; Zhang et al., 2023d; Han et al., 2023), and GeoAI (Zhao and Tang, 2017; Zhao et al., 2022; Zhang et al., 2023a, e). Previous research in image geolocalization could be primarily classified into two approaches: classification-based methods and retrieval-based methods. (1) Classification-based methods partition most regions of the Earth into multiple grid cells. Models are trained to classify each image into the correct cell (Clark et al., 2023b; Pramanick et al., 2022a; Müller-Budack et al., 2018; Seo et al., 2018a; Weyand et al., 2016b). The center coordinates of each cell are used as the predicted values. However, due to the limited number of cells, the granularity of the predicted values is coarse, preventing precise predictions. (2) Retrieval-based methods establish a database of geographic images with GPS coordinates. For a given input image, these methods retrieve the most similar image from the dataset and use its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018; Zhang et al., 2023b; Workman et al., 2015; Liu and Li, 2019; Zhou et al., 2024). However, constructing a comprehensive global-level image database is clearly impractical.
@@ 2.2. Geolocation Dataset
Existing geolocation datasets primarily originate from web-scraped or street-view images that have not been human-validated, raising concerns about their quality for effectively evaluating geolocation capabilities. For instance, datasets derived from web scraping, such as YFCC100M (Theiner et al., 2022b) and Im2GPS3K (Vo et al., 2017a), include a significant proportion of images depicting food, art, pets, and personal photographs. These types of images are often weakly localizable or entirely non-localizable (Theiner et al., 2022a). Street-view datasets also face limitations, such as restricted geographic coverage (Astruc et al., 2024); for example, (Mirowski et al., 2019)’s work includes data from only three cities in the United States. Furthermore, dataset collection processes often introduce biases. For instance, some commonly used platforms are inaccessible in certain countries, resulting in uneven geographic representation. Additionally, the difficulty of individual geolocation tasks varies widely within these datasets, but this aspect has not been comprehensively evaluated. For example, images taken at prominent landmarks are relatively easy to geolocate, while others offer no clear hints and are highly challenging (Astruc et al., 2024). These limitations hinder the ability of current datasets to provide a reliable benchmark for geolocation task.
Figure 3. Performance of game players of different levels in mainstream countries.
@@ 2.3. Large Vision Language Models
2.4. Vision Chain of Thought
The concept of Chain of Thought (CoT) was first introduced in the field of natural language processing, where explicit reasoning steps are modeled as intermediate representations to improve complex problem-solving (Wei et al., 2022; Fang et al., 2024). CoT has demonstrated significant improvements in tasks such as arithmetic reasoning, commonsense inference, and multi-step question answering (Dasgupta et al., 2022). Building on this foundation, recent research has adapted CoT reasoning to the realm of LVMs to enhance interpretability and robustness. For instance, (Wu et al., 2023) proposed a Description then Decision strategy to improve task decomposition and step-by-step processing. Furthermore, (Ge et al., 2023) advanced Vision CoT techniques to improve generalization in image classification tasks. In this work, we extend the CoT framework to the geolocation domain by introducing GeoCoT.
@@ 3.1. Geolocation Competition
Inspired by geoguessr.com, we developed a free geolocation game platform that tracks participants’ competition histories. Unlike most geolocation websites, including Geoguessr, which rely solely on samples from Google Street View, our platform integrates Baidu Maps and Gaode Maps to address coverage gaps in regions like mainland China, ensuring broader global accessibility. The platform offers various engaging competition modes to enhance user experience, such as team contests and solo matches. Each competition consists of multiple questions, and teams are assigned a “vitality score”. Users mark their predicted location on the map, and the evaluation is based on the ground truth’s surface distance from the predicted location. Larger errors result in greater deductions from the team’s vitality score. At the end of the match, the team with the higher vitality score wins. We also provide diverse game modes, including street views, natural landscapes, and iconic landmarks. Users can choose specific opponents or engage in random matches. To prevent cheating, external search engines are banned, and each round is time-limited. To ensure predictions are human-generated rather than machine-generated, users must register with a phone number, enabling tracking of individual activities. Using this platform, we collected GeoComp, a comprehensive dataset covering 1,000 days of user competition.
@@ 3.3. Human Player Performance
Table 2. Comparison of Precision, Recall and F1 scores in country-level and city-level geolocation. The scores are represented as follows: best, second, and third. Bold values indicate that our model achieves the best performance.
@@ 4. Geographic Chain of Thought
In this section, we introduce GeoCoT, a novel chain-of-thought prompting framework for graph-based and geolocation tasks. GeoCoT aims to improve LVMs’ performance on geographic reasoning tasks without relying on any additional prior information.
@@ 4.1. Rethinking Geolocation Task
The geolocation task has traditionally relied on classification-based (Clark et al., 2023b; Pramanick et al., 2022a; Weyand et al., 2016b) and retrieval-based methods (Zhu et al., 2022; Zhang et al., 2023b), as shown in Figure 4. There are also some works that combine these two paradigms together such as (Haas et al., 2024). While these approaches have advanced the field, they face significant limitations in precision and scalability, prompting a rethinking of the task. Classification-based methods divide the Earth into grid cells, predicting an image’s cell and using its center as the location. While simplifying the problem, this approach limits granularity, and increasing cell numbers adds computational complexity, hindering global use. Retrieval-based methods rely on geographically tagged databases but face scalability issues and poor performance in underrepresented regions, limiting effectiveness.
@@ 4.2. GeoCoT Deisgn
Specifically, our design of GeoCoT is inspired by the way humans approach geolocation tasks, progressing from broad to fine-grained analysis. When tasked with identifying a location from an image, humans naturally begin with macro-level observations, such as geographic features or climate, and progressively narrow their focus to country-specific, city-specific, and ultimately micro-level details. GeoCoT mimics this hierarchical reasoning to achieve accurate and interpretable geolocation predictions.
Concretely, GeoCoT operates in five sequential stages: 1. Continental or Climate Zone Identification. The process begins with identifying broad regions using natural features like mountains, vegetation, or soil, narrowing the scope to a continent or climate zone. 2. Country-Level Localization. Cultural markers, language on signs, and architectural styles are analyzed to refine predictions to the country level. 3. City-Level Refinement Using Infrastructure. Street elements, such as driving direction, bollards, and license plate colors, are used to locate specific cities or regions. 4. Landmark-Based Verification. Features like fire hydrants, guideposts, and street signs help validate and further refine the predicted location. 5. Fine-Grained Micro-Level Validation. Finally, subtle details such as sidewalk patterns and clothing styles confirm precise localization at a city or neighborhood level. Detailed prompts can be found in Appendix A. The answers generated at each of these stages are combined to produce the final predicted city, ensuring a comprehensive and step-by-step reasoning process guided by LVMs. By emulating human reasoning, GeoCoT surpasses traditional methods by avoiding coarse classification grids and exhaustive image databases, providing a scalable, interpretable, and accurate solution for geolocation tasks. It is important to note that GeoCoT does not require any concrete knowledge about the specific features of locations. Instead, it offers reasoning tutorials designed to help LVMs identify geographic clues by leveraging their existing knowledge.
@@ 5.1. Setting
We selected 500 geo-tagged locations with high inferential value from the dataset and employed a stratified sampling method across continents to ensure a balanced sample distribution. We define “high inferential value” as locations that present a moderate level of difficulty—challenging enough that many experienced participants can identify them, but not so simple that beginners can easily do so. Specifically, we selected 20 mainstream countries from six continents as representative samples and extracted tasks with an average player performance score of 3,000 out of 5,000 for annotation. Our GeoCoT is based on GPT-4o.
@@ 5.2. Baselines
We compare our model, GeoCoT, against several strong baselines representing the latest advancements in geolocation. General Open-Source VLMs: LLaVA-1.6 (Liu et al., 2023) utilizes a fully connected vision-language connector, effectively bridging visual inputs with linguistic features to deliver strong results in geolocation tasks. Llama-3.2-vision (Meta, 2024) demonstrates advanced multi-modal reasoning capabilities, making it a powerful open-source vision-language model. Qwen-VL (Bai et al., 2023), leveraging vast datasets of billions of image-text pairs, achieves robust performance in geolocation through its strong visual and spatial semantic understanding. Baselines Targeting Geolocation Tasks: PlaNet (Weyand et al., 2016c) partitions the Earth into cells and trains on millions of images for classification-based geolocalization. CPlaNet (Seo et al., 2018b) improves this by intersecting coarse partitions for fine-grained classes. Translocator (Pramanick et al., 2022b) uses a dual-branch transformer with RGB images and segmentation maps for robust planet-scale localization. GeoCLIP (Vivanco Cepeda et al., 2024), inspired by CLIP, aligns images with GPS coordinates using a retrieval-based approach to enhance geolocation. GeoReasoner (Li et al., 2024a) combines geospatial reasoning with visual-language alignment for state-of-the-art geolocation performance. Note that we report all baseline results on traditional benchmarks and partially on our dataset, as some models are not open-source, and others operate under different settings from our task. Closed-Source VLMs: GPT-4o (OpenAI, 2024) excels in vision reasoning tasks with its advanced multi-modal capabilities. Furthermore, GPT-4o with CoT (Wei et al., 2022) leverages chain-of-thought reasoning to improve zero-shot geolocation performance in complex scenarios.

### v2

@@ Abstract.
Geolocation, the task of identifying an image’s location, requires complex reasoning and is crucial for navigation, monitoring, and cultural preservation. However, current methods often produce coarse, imprecise, and non-interpretable localization. A major challenge lies in the quality and scale of existing geolocation datasets. These datasets are typically small-scale and automatically constructed, leading to noisy data and inconsistent task difficulty, with images that either reveal answers too easily or lack sufficient clues for reliable inference. To address these challenges, we introduce a comprehensive geolocation framework with three key components: GeoComp, a large-scale dataset; GeoCoT, a novel reasoning method; and GeoEval, an aspect-based metric designed to evaluate the correctness of the geolocation reasoning process. At the core of this framework is GeoComp (Geolocation Competition Dataset), a large-scale dataset collected from a geolocation game platform involving 740K users over two years. It comprises 25 million entries of metadata and 3.9 million geo-tagged locations spanning much of the globe, with each location annotated thousands to tens of thousands of times by human users. The dataset offers diverse difficulty levels for detailed analysis and highlights key gaps in current models. Building on this dataset, we propose Geographical Chain-of-Thought (GeoCoT), a multi-step reasoning framework designed to enhance the reasoning capabilities of Large Vision Models (LVMs) in geolocation tasks. GeoCoT improves performance by integrating contextual and spatial cues through a multi-step process that mimics human geolocation reasoning. Finally, we demonstrate that GeoCoT significantly boosts performance by up to 25% on classic geolocation metrics and by 9% in reasoning quality as measured by GeoEval:
@@
copyright: noneisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Google-WS-15k  (Clark et al., 2023a) 15k Global Map Service
GMCP (Zamir and Shah, 2014) 105K Local Map Service
StreetCLIP (Haas et al., 2023) 1M Unknown Map Service
Im2GPS  (Hays and Efros, 2008) 237 Local Web-Scraped
@@
Table 1. Comparison of Existing Geolocation Datasets and GeoComp. “Local” refers to city- or region-specific data, while “Global” spans multiple continents. Darker green shades indicate broader geographic coverage.
@@
To address the above obstacles, in this work, we leverage the contributions of hundreds of thousands of geolocation game enthusiasts who provide real user prediction annotations while playing the game. Specifically, we launched a free, public-benefit-oriented online geoguessing platform in June 2022, as shown in Figure 1(a). A screenshot of the platform’s GUI is provided in Appendix A. In each game, two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground-truth location. The images are sourced from Google Maps, Baidu Maps, Tencent Maps, and Gaode Maps. The platform offers multiple game modes, allowing users to either choose opponents or join random matchups. As of December 17, 2024, this platform has 740,468 users, 3,954,397 locations as unique geolocation tasks, and 25,355,174 human response records. We name the collected dataset GeoComp. This rich and valuable dataset of real human responses enables us to evaluate task difficulty and filter out unreasonable cases. For instance, some tasks are too easy, such as when the name of a shopping mall in a city is clearly visible in the image, enabling most users to answer correctly. On the other hand, some tasks are highly challenging, where only a few users spend considerable time before providing accurate answers. Additionally, there are unreasonable tasks that contain no identifiable hints, making them unsolvable for all users despite significant effort.
@@
Our work makes key contributions to geolocation. First, we present GeoComp, a large-scale, human-annotated geolocation dataset with over 3.9 million location images with corresponding location labels, and 25 million human player annotations, featuring diverse geographic regions, languages, and environmental contexts. These annotations identify high-difficulty geolocation cases and establish benchmarks to guide future advancements. Second, we introduce the Geographical Chain of Thought (GeoCoT) framework, a multi-step reasoning approach that improves geolocation accuracy by leveraging geographical cues like landmarks, environmental features, and spatial relationships. Finally, through comprehensive evaluations involving human assessments and LLM inferences, we show that GeoCoT improves predictive performance by up to 25% while enhancing interpretability.
Figure 1. (a) The gaming logic of our platform: Two players independently guess the location based on the same image and their own hints, with scores determined by the distance between their predictions and the ground truth location. (b) The global map shows spatial heterogeneity, with dense clusters in more urbanized regions like Europe and Asia, and sparse coverage in areas like Africa and Siberia. (c) The pie chart highlights the proportional geo-tagged locations distribution, led by North America and Asia. (d) Unlike previous datasets like OSV-5M, where a single country (e.g., America) dominates 25% of the data, our dataset is balanced at country level.
@@ 2.1. Image Geolocation Task
Image geolocation refers to determining the corresponding location of a given image, a crucial task in computer vision (Zhu et al., 2023b, a, 2024a), spatial data mining (Zhao et al., 2017, 2016; Zhang et al., 2023d; Han et al., 2023), and GeoAI (Zhao and Tang, 2017; Zhao et al., 2022; Zhang et al., 2023a, e). Previous research in image geolocalization could be primarily classified into two approaches: classification-based methods and retrieval-based methods. (1) Classification-based methods partition most regions of the Earth into multiple grid cells. Models are trained to classify each image into the correct cell (Clark et al., 2023b; Pramanick et al., 2022; Müller-Budack et al., 2018; Seo et al., 2018; Weyand et al., 2016b). The center coordinates of each cell are used as the predicted values. However, due to the limited number of cells, the granularity of the predicted values is coarse, preventing precise predictions. (2) Retrieval-based methods establish a database of geographic images with GPS coordinates. For a given input image, these methods retrieve the most similar image from the dataset and use its coordinates as the predicted location (Zhu et al., 2022; Müller-Budack et al., 2018; Zhang et al., 2023b; Workman et al., 2015; Liu and Li, 2019; Zhou et al., 2024). However, constructing a comprehensive global-level image database is clearly impractical.
@@ 2.2. Geolocation Dataset
Existing geolocation datasets primarily originate from web-scraped or street-view images that have not been human-validated, raising concerns about their quality for effectively evaluating geolocation capabilities. For instance, datasets derived from web scraping, such as YFCC100M (Theiner et al., 2022b) and Im2GPS3K (Vo et al., 2017a), include a significant proportion of images depicting food, art, pets, and personal photographs. These types of images are often weakly localizable or entirely non-localizable (Theiner et al., 2022a). Street-view datasets also face limitations, such as restricted geographic coverage (Astruc et al., 2024); for example, (Mirowski et al., 2019)’s work includes data from only three cities in the United States. Furthermore, dataset collection processes often introduce biases. For instance, some commonly used platforms are inaccessible in certain countries, resulting in uneven geographic representation. Additionally, the difficulty of individual geolocation tasks varies widely within these datasets, but this aspect has not been comprehensively evaluated. For example, images taken at prominent landmarks are relatively easy to geolocate, while others offer no clear hints and are highly challenging (Astruc et al., 2024). These limitations undermine the reliability of current geolocation benchmarks.
Figure 2. Performance of game players of different levels in mainstream countries. Experts are defined as the top 15% in performance scores, while beginners are those in the bottom 15%.
@@ 3.1. Geolocation Competition
Inspired by geoguessr website, we developed a free geolocation game platform that tracks participants’ competition histories. Unlike most geolocation websites, including Geoguessr, which rely solely on samples from Google Street View, our platform integrates Baidu Maps, Tencent Maps, and Gaode Maps to address coverage gaps in regions like mainland China, ensuring broader global accessibility. Users can choose specific opponents or engage in random matches. Each match consists of multiple questions, and each user is initially assigned a “vitality score.” Users mark their predicted location on a map, and the system evaluates accuracy based on the surface distance between the predicted point and the ground truth. Larger errors result in greater deductions from the user’s vitality score. The user with the higher vitality score at the end of the match is declared the winner. To ensure predictions are human-generated rather than machine-generated, users must register with a phone number, enabling tracking of individual activities. Using this platform, we collected GeoComp, a comprehensive dataset covering 1,000 days of user competition.
@@ 3.3. Human Player Performance
Table 2. Comparison of Precision, Recall and F1 scores in country-level and city-level geolocation. The scores are represented as follows: best, second, and third. Numbers in bold mean that the improvement to the best baseline is statistically significant (a two-tailed paired t-test with p-value <0.01).
@@ 4. Geographic Chain of Thought
In this section, we introduce GeoCoT, a novel chain-of-thought prompting framework for graph-based and geolocation tasks. Unlike standard CoT prompting which performs generic step-by-step reasoning, GeoCoT introduces a domain-specific, hierarchically structured reasoning process that mimics how humans localize geographic information from broad regions to fine-grained details.
@@ 4.1. Rethinking Geolocation Task
As discussed in § 2.1, the geolocation task has traditionally relied on classification-based (Clark et al., 2023b; Pramanick et al., 2022; Weyand et al., 2016b) and retrieval-based methods (Zhu et al., 2022; Zhang et al., 2023b), as shown in Figure 3. While these approaches have advanced the field, they face significant limitations in precision and scalability, prompting a rethinking of the task.
@@ 4.2. GeoCoT Deisgn
Our design of GeoCoT is inspired by how humans intuitively approach geolocation—progressing from broad to fine-grained analysis. Rather than relying on generic step-by-step reasoning like standard CoT prompting, GeoCoT mimics the human process: starting with macro-level cues (e.g., climate, terrain), then narrowing down to country, city, and finally micro-level details to guide the model through interpretable geographic reasoning.
Concretely, GeoCoT operates in five sequential stages: 1. Continental or Climate Zone Identification. The process begins with identifying broad regions using natural features like mountains, vegetation, or soil, narrowing the scope to a continent or climate zone. 2. Country-Level Localization. Cultural markers, language on signs, and architectural styles are analyzed to refine predictions to the country level. 3. City-Level Refinement Using Infrastructure. Street elements, such as driving direction, bollards, and license plate colors, are used to locate specific cities or regions. 4. Landmark-Based Verification. Features like fire hydrants, guideposts, and street signs help validate and further refine the predicted location. 5. Fine-Grained Micro-Level Validation. Finally, subtle details such as sidewalk patterns and clothing styles confirm precise localization at a city or neighborhood level. These five reasoning steps are formulated as a single, structured prompt and jointly fed into the LVM, which directly generates the final predicted location. Detailed prompts can be found in Appendix B.
It is important to note that GeoCoT does not require any concrete knowledge about the specific features of locations. Instead, it offers reasoning tutorials designed to help LVMs identify geographic clues by leveraging their existing knowledge.
@@ 5.1. Setting
We selected 500 geo-tagged locations with high inferential value from the dataset to serve as a test set, using a stratified sampling method across continents to ensure balanced geographic distribution. This number is larger than in previous works (Liu et al., 2023b; Guan et al., 2024), which typically include only a few dozen case studies to examine the reasoning process. We define ”high inferential value” as locations with moderate difficulty—challenging enough to be correctly identified by experienced participants, but not so easy that beginners can do so effortlessly. Specifically, we selected 20 mainstream countries across six continents as representative samples and extracted tasks with an average player score of around 3,000 out of 5,000 for annotation. This test set has been publicly released on GitHub. Our GeoCoT framework is implemented using GPT-4o.
@@ 5.2. Baselines
We compare our model, GeoCoT, against several strong baselines representing the latest advancements in geolocation on our dataset. General Open-Source VLMs: LLaVA-1.6 (Liu et al., 2023a) utilizes a fully connected vision-language connector, effectively bridging visual inputs with linguistic features to deliver strong results in geolocation tasks. Llama-3.2-vision (Meta, 2024) demonstrates advanced multi-modal reasoning capabilities, making it a powerful open-source vision-language model. Qwen-VL (Bai et al., 2023), leveraging vast datasets of billions of image-text pairs, achieves robust performance in geolocation through its strong visual and spatial semantic understanding. Baselines Targeting Geolocation Tasks: GeoCLIP (Vivanco Cepeda et al., 2024), inspired by CLIP, aligns images with GPS coordinates using a retrieval-based approach to enhance geolocation. GeoReasoner (Li et al., 2024a) combines geospatial reasoning with visual-language alignment for state-of-the-art geolocation performance. Closed-Source VLMs: GPT-4o (OpenAI, 2024) excels in vision reasoning tasks with its advanced multi-modal capabilities. Furthermore, GPT-4o(CoT), following the setting of cot-zero-shot (Wei et al., 2022), leverages chain-of-thought reasoning to improve performance in complex scenarios. All models are evaluated using the same input format and test set to ensure a fair comparison.